OpenStudy (anonymous):

Find the slope of(-4,6) (3,2) and write an equation y-y1=m(x-x1)

OpenStudy (anonymous):

squeaky try rmbr the Formula!

OpenStudy (angela210793):

slope=(y2-y1)/(x2-x1)=-4/7 y-6=-4/7(x+4)

OpenStudy (anonymous):

\[Slope = (y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

DOn't See the solution.... TRy Grasp the Concepts Behind it.

OpenStudy (anonymous):

TRY SOLVIng with points (9,0) and (4,4) .. !